OpenStudy (cookie_2046):

6x-2=22 what is x = to?

OpenStudy (kaylardz0405):

22 divided by 6

OpenStudy (insa):

6x=22+2 6x=24 x=24/6

OpenStudy (samigupta8):

4

OpenStudy (insa):

x=4

OpenStudy (cookie_2046):

Thank u soo much

OpenStudy (insa):

NP :)